Six litres of 0.4M 3-(N-morpholino) propane sulfonic acid (MOPS) buffer of pH 7.8 are required. Available in the laboratory are 10M HCl and 10M NaOH and the zwitterionic form of MOPS (M.Wt = 210.3). The pKa for the MOPS zwitter ion is 7.2. Describe, showing the relevant calculations, the procedure for preparing the buffer
1
Expert's answer
2020-09-16T05:10:47-0400

pH = pKa ​+ log (CBA / CHA)

7.8 = 7.2 + log (CBA / CHA)

log (CBA / CHA) = 0.6


So:

CBA / CHA = 3.98

and CBA + CHA = 0.4


Solving this system of equations, we have:

CHA ​= 0.08M

CBA ​= 0.4 − CHA ​= 0.32M


"n(HA)=C_{HA}*V_{sol}=0.08M*6L=0.48mol"

"n(BA)=C_{BA}*V_{sol}=0.32M*6L=1.92mol"


"n(HA)+n(BA)=0.48mol+1.92mol=2.40mol"

Now we take 2.40 mol * 209.3 g/mol = 502 g of MOPS 

and 1.92 mol/10M = 0.192L of 10M solution of NaOH (react in a 1 to 1 molar relation):

C7H15NO4S + NaOH = C7H14NO4SNa + H2O

After the reaction is complete, you dissolve this mixture up to 6L in a special flask with distilled water.